OpenStudy (anonymous):

3x - 2 / 12 - 1/6 = 1/6? Could someone show me all the steps to do this problem?

OpenStudy (anonymous):

\[\frac{ 3x -2 }{ 12 } - \frac{ 1 }{ 6 } = \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

\[\frac{ 3x-2 }{ 12 }=\frac{ 2 }{ 6 }\] \[3x-2=2\frac{ 12 }{ 6 }\] \[3x-2=4\] \[3x=6\] \[x=2\]

OpenStudy (anonymous):

Thank you, could you help me with some more?

OpenStudy (anonymous):

I can